# Optimistic Safety for Linearly-Constrained Online Convex Optimization

### v1

@@ Appendix D Example OCO Algorithms
In this section, we give examples of algorithms that can be used for the OCO algorithm $\mathcal{A}$ in OSOCO, i.e. those that admit the type of regret bound specified in Definition 3.2.
D.1 Problem setting
@@
We consider a setting where at each round $t$, the player chooses an action $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and, then the player observes the cost function $f_{t}$ and suffers the cost $f_{t}(x_{t})$. The game can be terminated at any time and therefore the final round $T$ is unknown to the player. In the algorithms below, we use the flag $\mathsf{alive}$ to indicate that the game has not been terminated. Also, $\mathcal{X}$ is not necessarily convex. We assume that all $f_{t}$ are differentiable and $\|\nabla f_{t}(x)\|\leq G$ for all $x\in\mathcal{X}$. We also assume that $\|x\|\leq D/2$ for all $x\in\mathcal{X}$. We also take the adversary to be oblivious such that the cost functions are fixed ahead of time.
D.2 Example Algorithm 1: Anytime Hedge over a Finite Cover
@@
The simplest algorithm to handle this problem is Hedge played over a finite cover with the doubling trick. We give pseudocode for this algorithm in Algorithm 3. This is a well-known extension of the Hedge algorithm Freund & Schapire (1997) that has been noted by works such as Dani et al. (2007), Maillard & Munos (2010) and Krichene et al. (2015).
@@
$\begin{split}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\min_{m\in M_{j}}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(z_{m})&=GD\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[\ell_{t}(x_{t})]-GD\min_{m\in M_{j}}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\ell_{t}(z_{m})\\ &\leq GD\sqrt{4\tau_{j}\log(M_{j})},\end{split}$ (8)
@@
$\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x^{\star}_{j})$ $\leq\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(z_{m^{\star}_{j}})+\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(z_{m^{\star}_{j}})-\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x^{\star}_{j})$
@@
Furthermore, it is known (e.g. Example 27.1 in Shalev-Shwartz & Ben-David (2014)) that we can construct a grid for the net to get $M_{j}\leq(2D\sqrt{d}/\Delta_{j})^{d}$ for desired $\Delta_{j}$. Therefore, choosing $\Delta_{j}=1/(G\sqrt{\tau_{j}})$ will need $M_{j}\leq(2DG\sqrt{d\tau_{j}})^{d}$ points. Also, note that $N$ is the smallest integer such that $\sum_{j=0}^{N}2^{j}=2^{N+1}-1\geq T$ and therefore $N=\lceil\log_{2}(T+1)-1\rceil\leq\log_{2}(T+1)$. Therefore, it holds that
@@
D.3 Example Algorithm 2: Anytime Follow the Perturbed Leader
Another algorithm that can be used for this setting is Follow the Perturbed Leader (FTPL) Kalai & Vempala (2005) which has been popular for a variety of OCO settings. We use the version from Suggala & Netrapalli (2020) with an exact optimization oracle. We use the version and analysis from Suggala & Netrapalli (2020) because they study it in the nonconvex setting, which is directly applicable to our setting because the action set is nonconvex. We give the FTPL algorithm with the doubling trick in Algorithm 4.
@@ D.3 Example Algorithm 2: Anytime Follow the Perturbed Leader
We will use the analysis from Suggala & Netrapalli (2020) to bound the regret within each phase. First note that Suggala & Netrapalli (2020) assumes that $\|x-y\|_{\infty}\leq D^{\prime}$ for all $x,y\in\mathcal{X}$ and that $f_{t}$ are $L^{\prime}$-Lipschitz w.r.t. the 1-norm. In our setting, it holds that for all $x,y\in\mathcal{X}$ that
@@
Therefore, we can take $D^{\prime}=D$ and $L^{\prime}=G$. Also, Suggala & Netrapalli (2020) assume access to a $(\alpha,\beta)$ inexact oracle, but we assume that this oracle is exact and therefore take $\alpha,\beta=0$. Then, since the distribution over actions chosen by FTPL in each round (denoted by $p_{t}$) is independent of past randomness when the adversary is oblivious, we can apply Theorem 1 in Suggala & Netrapalli (2020) to get that
@@
Appendix E Numerical Experiments
In this appendix, we provide further details on the numerical experiments from Section 7 and additional results. Specifically, we discuss the linear-cost setting in Section E.1, the quadratic cost setting in Section E.2.
E.1 Safe Online Linear Programming
First, we give details on the setting in Figure 0(a) and then give additional results.
Algorithm Parameters for Figure 0(a)
First, we give the problem parameters defined by the assumptions:
  • $D=2$
  • $S=\sqrt{2}$
  • $G=\sqrt{2}$
  • $\rho=0.01$
  • $b_{\mathrm{min}}=0.6$
Then, the algorithm parameters for SO-PGD are chosen as suggested by Chaudhary & Kalathil (2022):
  • $\gamma=b_{\mathrm{min}}/S$
  • $\eta=D/(G\sqrt{T})$
  • $T_{0}=T^{2/3}$
  • $\rho=0.01$
  • $\lambda=1$
  • $\delta=0.01$
Also, OSOCO is run with HedgeDescent for $\mathcal{A}$ and the algorithm parameters are chosen as follows:
  • $\eta_{t}=D/(G\sqrt{t})$ (in HedgeDescent)
  • $\zeta_{t}=\sqrt{4\log(2d)}/(GD\sqrt{t})$ (in HedgeDescent)
  • $\lambda=1$
  • $\delta=0.01$
Note that this choice of $\lambda$ differs from the one specified in Theorem 4.1. However, this choice will only change the regret guarantees by constant factors.
Additional Results
Figure 2: The performance of OGD Zinkevich (2003) with full constraint information and OSOCO in a “hard” setting. The line shows the average regret values over $60$ trials for $T\in\{1000,2000,...,100000\}$ and the shading shows $\pm$ 1 standard deviation.
We also give additional results in a different setting. In particular, we consider a setting that is challenging for online algorithms in general, rather than just safe algorithms. This differs from the settings in Section 7, which are intended to demonstrate the performance of OSOCO in a challenging safe learning setting. The results for OGD Zinkevich (2003) (with full constraint information) and OSOCO are given in Figure 2, where the line shows the regret with $T\in\{1000,2000,...,100000\}$ and is averaged over $60$ trials and the shading shows ±$\pm 1$ standard deviation. The setting is defined as:
  • $d=2$
  • $\mathcal{X}=\sqrt{2}\mathbb{B}$
  • $A=[I\ -I]^{\top}$
  • $\theta_{t}\sim\{-1,1\}^{d}$
  • $\epsilon_{t}\sim\mathcal{U}(0,\rho I)$
  • $\rho=1$
  • $b=\mathbf{1}$
  • $D=2\sqrt{2}$
  • $S=\sqrt{2}$
  • $G=\sqrt{2}$
  • $b_{\mathrm{min}}=1$
Note that this setting is very challenging for both online algorithms in general and safe algorithm in particular. This is evident from the fact that there is a lot of noise on the constraint feedback ($\epsilon_{t}$ has standard deviation $1$), and yet OGD and OSOCO perform similarily.
In this setting, HedgeDescent is used for OSOCO with the following algorithm parameters:
  • $\eta_{t}=D/(G\sqrt{t})$ (in HedgeDescent)
  • $\zeta_{t}=\sqrt{4\log(2d)}/(GD\sqrt{t})$ (in HedgeDescent)
  • $\lambda=1$
  • $\delta=0.01$
Also, $\eta_{t}=D/(G\sqrt{t})$ is used for OGD Zinkevich (2003).
E.2 Safe Online Quadratic Programming
First, we give the algorithm parameters used for Figure 0(b) and then give some additional results.
Algorithm Parameters for Figure 0(b)
First, we give the problem parameters defined by the assumptions:
  • $D=2$
  • $S=\sqrt{2}$
  • $G=4\sqrt{2}+4$
  • $\rho=0.01$
  • $b_{\mathrm{min}}=0.5$
Then, the algorithm parameters for SO-PGD are chosen as suggested by Chaudhary & Kalathil (2022):
  • $\gamma=b_{\mathrm{min}}/S$
  • $\eta=D/(G\sqrt{T})$
  • $T_{0}=T^{2/3}$
  • $\rho=0.01$
  • $\lambda=1$
  • $\delta=0.01$
Also, OSOCO is run with HedgeDescent for $\mathcal{A}$ and the algorithm parameters are chosen as follows:
  • $\eta_{t}=D/(G\sqrt{t})$ (in HedgeDescent)
  • $\zeta_{t}=\sqrt{4\log(2d)}/(GD\sqrt{t})$ (in HedgeDescent)
  • $\lambda=1$
  • $\delta=0.01$
Additional Results
Figure 3: The performance of the algorithm in Yu et al. (2017) (abbreviated DPP in reference to “drift-plus penalty”) with full constraint feedback and OSOCO. The points show the average regret values over $30$ trials and the error bars show $\pm$ 1 standard deviation.
Figure 4: The instantaneous violation of the algorithm in Yu et al. (2017) (labeled DPP) with full constraint feedback and OSOCO. The line shows the average regret values over $30$ trials and the shading shows $\pm$ 1 standard deviation.
We also give results comparing OSOCO with the algorithm from Yu et al. (2017). The results of this are shown in Figure 3, where the algorithm from Yu et al. (2017) is referred to as DPP in reference to “drift-plus-penalty”, the name for their analysis technique. The plotted (cumulative) violation is defined as $V_{T}=\sum_{t=1}^{T}(y_{t}-b)$. In this case, we find that the regret of OSOCO is closer to constant $\sqrt{T\log(T)}$ and therefore use that normalization factor. Note that the violation for OSOCO is not shown because it is negative for all trials. We also plot the instantaneous violation for a fixed horizon in Figure 4. The setting is defined as:
@@ E.2 Safe Online Quadratic Programming
In this setting, HedgeDescent is used for OSOCO with the following algorithm parameters:
@@
  • $\mu=b_{\mathrm{min}}/T$ (in the sense of Proposition 5.1)
Also, DPP uses the following parameters as recommended in Yu et al. (2017):
  • $\alpha=T$

### v3

@@ Appendix F Example OCO Algorithms
In this section, we give examples of algorithms that can be used for the OCO algorithm $\mathcal{A}$ in OSOCO, i.e. those that admit the type of regret bound specified in Definition 1.
F.1 Problem setting
@@
We consider a setting where at each round $t$, the player chooses an action $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and an adversary simultaneously chooses a cost function $f_{t}$ given the actions played in previous rounds. The player then observes $f_{t}$ and suffers the cost $f_{t}(x_{t})$. The game can be terminated at any time and therefore the final round $T$ is unknown to the player. In the algorithms below, we use the flag $\mathsf{alive}$ to indicate that the game has not been terminated. Also, $\mathcal{X}$ is not necessarily convex. We assume that all $f_{t}$ are differentiable and $\|\nabla f_{t}(x)\|\leq G$ for all $x\in\mathcal{X}$. We also assume that $\|x\|\leq D/2$ for all $x\in\mathcal{X}$.
F.2 Example Algorithm 1: Anytime Hedge over a Finite Cover
@@
The simplest algorithm to handle this problem is Hedge played over a finite cover with the doubling trick. We give pseudocode for this algorithm in Algorithm 3. This is a well-known extension of the Hedge algorithm Freund and Schapire (1997) that has been noted by works such as Dani et al. (2007), Maillard and Munos (2010) and Krichene et al. (2015).
@@
$\begin{split}&\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\min_{m\in M_{j}}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(z_{m})\\ &=GD\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[\ell_{t}(x_{t})]-GD\min_{m\in M_{j}}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\ell_{t}(z_{m})\\ &\leq GD\sqrt{4\tau_{j}\log(M_{j})},\end{split}$ (8)
@@
$\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x^{\star}_{j})$
$\leq\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(z_{m^{\star}_{j}})+\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(z_{m^{\star}_{j}})-\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x^{\star}_{j})$
@@
Furthermore, it is known (e.g. Example 27.1 in Shalev-Shwartz and Ben-David (2014)) that we can construct a grid for the net to get $M_{j}\leq(2D\sqrt{d}/\Delta_{j})^{d}$ for desired $\Delta_{j}$. Therefore, choosing $\Delta_{j}=1/(G\sqrt{\tau_{j}})$ will need $M_{j}\leq(2DG\sqrt{d\tau_{j}})^{d}$ points. Also, note that $N$ is the smallest integer such that $\sum_{j=0}^{N}2^{j}=2^{N+1}-1\geq T$ and therefore $N=\lceil\log_{2}(T+1)-1\rceil\leq\log_{2}(T+1)$. Therefore, it holds that
@@
F.3 Example Algorithm 2: Anytime Follow the Perturbed Leader
Another algorithm that can be used for this setting is Follow the Perturbed Leader (FTPL) Kalai and Vempala (2005) which has been popular for a variety of OCO settings. We use the version from Suggala and Netrapalli (2020) with an exact optimization oracle. We use the version and analysis from Suggala and Netrapalli (2020) because they study it in the nonconvex setting, which is directly applicable to our setting because the action set is nonconvex. We give the FTPL algorithm with the doubling trick in Algorithm 4.
@@ F.3 Example Algorithm 2: Anytime Follow the Perturbed Leader
We will use the analysis from Suggala and Netrapalli (2020) to bound the regret within each phase. First note that Suggala and Netrapalli (2020) assumes that $\|x-y\|_{\infty}\leq D^{\prime}$ for all $x,y\in\mathcal{X}$ and that $f_{t}$ are $L^{\prime}$-Lipschitz w.r.t. the 1-norm. In our setting, it holds that for all $x,y\in\mathcal{X}$ that
@@
Therefore, we can take $D^{\prime}=D$ and $L^{\prime}=G$. Also, Suggala and Netrapalli (2020) assume access to a $(\alpha,\beta)$ inexact oracle, but we assume that this oracle is exact and therefore take $\alpha,\beta=0$. Then, since the regret guarantees of FTPL hold against a non-oblivious adversary (see Theorem 8 in Suggala and Netrapalli (2020)), we can apply the regret guarantees of FTPL (Theorem 1 in Suggala and Netrapalli (2020)) to get that,
@@
Appendix G Details on Numerical Experiments
In this appendix, we provide details on the numerical experiments from Section 7. We consider a $2$-dimensional setting with quadratic costs of the form $f_{t}(x)=3\|x-v_{t}\|^{2}$. We choose $\mathcal{X}=\mathbb{B}$, $A=[-1\ -1]$, $b=0.8$ and sample $v_{t}\sim[-1,0]^{d}$ and $\epsilon_{t}\sim\mathcal{N}(0,\sigma I)$ where $\sigma=0.01$. In this setting, we consider OSOCO configured for high-probability (OSOCO-H) and OSOCO configured for static expectation (OSOCO-E) alongside the algorithm from Yu et al. (2017) (which we call DPP for "drift-plus-penalty") and SO-PGD from Chaudhary and Kalathil (2022). For DPP, we consider both the nominal choice of algorithm parameters as well as a conservative choice of algorithm parameters that are chosen to keep the violation very small (referred to as DPP(cons.)). The choice of algorithm parameters for all algorithms is given in Table 1.
We study the regret, the stochastic constraint violation and the static constraint violation as reported in Figure 2(a), Figure 2(b) and Figure 2(c), respectively. All figures show the results for $30$ trials, where the average is indicated by a point and the standard deviation is indicated by an errorbar or shading. These results indicate that the nominal DPP algorithm has lower regret than any other algorithm, although it also incurs significant violation of both static and stochastic constraints. When the algorithm parameters for DPP are chosen to ensure very little violation, the regret becomes substantial. The algorithms with the next lowest regret are OSOCO-T and OSOCO-S, which do not incur any stochastic constraint violation (on average) or static constraint violation.
Next, we give the details on the setting, and then give the algorithm parameters for each of the algorithms in Table 1. Note that, unless otherwise specified, these are chosen as recommended in each particular paper. In total, the experiments took about $30$ hours of CPU compute time.
Setting Details
@@ Appendix G Details on Numerical Experiments
Table 1: Choice of algorithm parameters for simulations.
[…]
@@
$\beta_{t}=\rho\sqrt{d\log\left(\frac{1+(t-1)D^{2}/\lambda}{\delta/n}\right)}+\sqrt{\lambda}S$
$\kappa=b_{\mathrm{min}}/T$
H $\mathcal{A}=\mathrm{HedgeDescent}$
$\eta_{t}=D/(G\sqrt{t})$ (in HedgeDescent)
$\zeta_{t}=\sqrt{4\log(2d)}/(GD\sqrt{t})$ (in HedgeDescent)
$\lambda=\max(1,D^{2})$
$\delta=0.01$
$\beta_{t}=\rho\sqrt{d\log\left(\frac{1+(t-1)D^{2}/\lambda}{\delta/n}\right)}+\sqrt{\lambda}S$
$\kappa=0$
DPP - $\alpha=T$
@@ H $\mathcal{A}=\mathrm{HedgeDescent}$
(cons.) $\alpha=T$
$V=0.01\sqrt{T}$
SO-PGD - $\lambda=1$
$\delta=0.01$
$\eta=D/(G\sqrt{T})$
$T_{0}=T^{2/3}$
[…]